# Gravitational wave probes of particle dark matter: a review

### v3

@@ XI Conclusions
There is no shortage of observable signatures that DM could leave on GW sources when the particle, mass and couplings of DM the standard model are varied. We do not assume to have covered every single DM model in this work, but we hope to have provided an overview for the interested reader that wishes to pursue more details about different ways that GWs can be used to probe the existence of particle DM.
@@
In this work, we have explored (1) how ultralight dark matter could couple to GW interferometers, (2) how macroscopic DM could transit through the interferometers, (3) how boson cloud systems formed through a superradiance process around rotating BHs could emit GWs by annihilation, (4) how soliton DM could induce changes in GWs arising from binary systems and isolated neutron stars, (5) how WIMP DM could collect around celestial objects and induce them to collapse into BHs detectable via GW emission, (6) how environmental effects of DM spikes around BHs could affect GW signals in future detectors, and (7) how GW observations and non-observations can be used to probe atomic models of DM. Some of these types for DM have been extensively constrained through analysis of terrestrial GW detector data; others will require the advent of space-based and next-generation ground-based GW interferometers to obtain insight into these types of DM.
@@
Despite the “WIMP miracle” announced decades ago, WIMPs have not yet been detected, and are only one of numerous possible ways that DM could exist in the Universe. The lack of detection of WIMPs over the years has opened up both the experimental and theoretical playgrounds for new probes of DM, one of which is through its impact on GW signals, or as a signal in interferometer data itself. Moreover, it has been assumed in this work, and almost all those cited, that DM is the way to solve the puzzling cosmological observations on a variety of scales. However, modified theories of gravity are real contenders to solve this problem as well, though their community is a lot smaller and not quite as popular as the DM one. These two communities – DM and modified gravity – act entirely independently, choosing to break, respectively, one of the two assumptions of “classical” physics: (1) the matter in the universe is mostly baryonic and luminous, and (2) general relativity and, in the limit of slow-moving objects, Newtonian theory, are correct Martens and Lehmkuhl (2020). It may be that these assumptions need to be broken simultaneously, or that these two distinct fields have more in common than it appears, as argued in Martens and Lehmkuhl (2020).
@@
We only briefly mentioned PBHs in this review, which have been reviewed extensively elsewhere Green and Kavanagh (2021); Miller (2024). PBHs could also comprise a fraction of or the totality of DM, and recent observations of binary black holes by the LIGO, Virgo and KAGRA with low spins whose merging rates are consistent with PBHs have reignited interested in this hypothesis Bird et al. (2016); Clesse and García-Bellido (2017); Sasaki et al. (2016); Croon et al. (2023). While analyses of PBHs typically receive criticism for using monochromatic mass functions to place constraints, all of the results presented here implicitly assume that DM is only composed of particles that are being constrained [^7] The exception to this statement is superradiance probes of BHs– these bosons need not be DM.. It could be that multiple forms of DM co-exist, e.g. in the atomic DM models, or that there is a whole dark sector.
@@
Nevertheless, GWs provide a new window to probe the existence of DM in a variety of ways. Most of the actual analyses performed on GW data do not assume particular models for DM, and only constrain them later when one interprets null (or actual) results in terms of a DM model. GW searches are therefore powerful, relatively model-agnostic probes of DM that permit the detection of a variety of DM signatures.
@@ Acknowledgments
We thank William East, Alexandre Gottel, Dana Jones, Jun’ya Kume, Sachiko Kuroyanagi, Ling Sun, and Yue Zhao for reviewing this manuscript.

### v4

@@ XI Conclusions
There is no shortage of observable signatures that DM could leave on GW sources when the particle, mass and couplings of DM the standard model are varied. We do not assume to have covered every single DM model in this work, but we hope to have provided an overview for the interested reader that wishes to pursue more details about different ways that GWs can be used to probe particle DM.
@@
In this work, we have explored (1) how ultralight dark matter could couple to GW interferometers, (2) how macroscopic DM can transit through the interferometers, (3) how boson cloud systems formed through a superradiance process around rotating BHs can emit GWs by annihilation, (4) how soliton DM can induce changes in GWs arising from binary systems and isolated neutron stars, (5) how WIMP DM can collect around celestial objects and induce them to collapse into BHs detectable via GW emission, (6) how environmental effects of DM spikes around BHs can affect GW signals in future detectors, and (7) how GW observations and non-observations of binary black holes can be used to probe atomic models of DM. Some of these types for DM have been extensively constrained through analysis of terrestrial GW detector data; others will require the advent of space-based and next-generation ground-based GW interferometers to obtain insight into these types of DM.
@@
Despite the “WIMP miracle” announced decades ago, WIMPs have not yet been detected, and are only one of numerous possible ways that DM could exist in the Universe. The lack of detection of WIMPs over the years has opened up both the experimental and theoretical playgrounds for new probes of DM, one of which is through its impact on GW signals, or as a signal in interferometer data itself. Moreover, it has been assumed in this work, and almost all those cited, that DM is the way to solve the puzzling cosmological observations on a variety of scales. However, modified theories of gravity are real contenders to solve this problem as well. These two communities – DM and modified gravity – act entirely independently, choosing to break, respectively, one of the two assumptions of “classical” physics: (1) the matter in the universe is mostly baryonic and luminous, and (2) general relativity and, in the limit of slow-moving objects, Newtonian theory, are correct Martens and Lehmkuhl (2020). It may be that these assumptions need to be broken simultaneously, or that these two distinct fields have more in common than it appears, as argued in Martens and Lehmkuhl (2020).
@@
We only briefly mentioned PBHs in this review, which have been reviewed extensively elsewhere Green and Kavanagh (2021); Miller (2024a). PBHs could also comprise a fraction of or the totality of DM, and recent observations of binary black holes by the LIGO, Virgo and KAGRA with low spins whose merging rates are consistent with PBHs have reignited interested in this hypothesis Bird et al. (2016); Clesse and García-Bellido (2017); Sasaki et al. (2016); Croon et al. (2023). While analyses of PBHs typically receive criticism for using monochromatic mass functions to place constraints, all of the results presented here implicitly assume that DM is only composed of particles that are being constrained [^13] The exception to this statement is superradiance probes of BHs– these bosons need not be DM.. It could be that multiple forms of DM co-exist, e.g. in the atomic DM models, or that there is a whole dark sector.
@@
Nevertheless, GWs provide a new window to probe the origins of DM in a variety of ways. Most of the actual analyses performed on GW data do not assume particular models for DM, and only constrain them later when one interprets results in terms of a DM model. GW searches are therefore powerful, relatively model-agnostic probes of DM that permit the detection of a variety of DM signatures.
@@ Acknowledgments
We thank William East, Alexandre Gottel, Dana Jones, Jun’ya Kume, Sachiko Kuroyanagi, Ling Sun, and Yue Zhao for reviewing this manuscript. We also thank the three anonymous referees that provided valuable feedback on this review article.